Bullseye Corp. does not pay dividends. As an analyst for the company, you are trying to assess the stock price and realize you cannot use the dividend discount model because of the lack of dividend payment.

For the coming year, you estimate the following based upon past performance and current conditions:

 EBIT of $500,000
 Depreciation of $250,000
 Tax rate of 21%
 Change in NWC of $100,000
 Capital Expenditures of $350,000
 Market Value of Debt of $3,000,000
 Excess Cash of $250,000
 100,000 shares outstanding
 W ACC=10.9%
 Expected growth in free cash flows of:
o Year 2 and Year 3: 10%
o Year 4 and beyond: 5%

What is the your estimate for Bullseye’s stock price?
Answer #1

Expected FCF next year = EBIT*(1-Tax rate) + Depreciation - Change in NWC - Capital expenditure

= 500,000*(1-21%) + 250,000 - 100,000 - 350,000 = 195,000

Valuation table:

Stock price = $8.52 per share